When the chemical equation SO2+Cl2–>SOCl2+Cl2O is balanced using the smallest set of whole number coefficients, what is the coefficent for SOCl2?
1
Expert's answer
2020-11-25T02:20:25-0500

Q146406

When the chemical equation SO2 + Cl2 –> SOCl2 + Cl2O is balanced using the smallest set of whole number coefficients, what is the coefficient for SOCl2 ?


Solution:


We will first count the total number of atoms on both reactant and product side.


SO2 + Cl2 –> SOCl2 + Cl2O

Reactant / Product


S 1 / 1


O 2 / 1 + 1 = 2


Cl 2 / 2 + 2 = 4


We can see that all the elements are balanced in this reaction except Cl.

So we will balance the Cl on both the side. 


Reactant Cl2 contains only Cl. So we can balance the Cl by writing proper coefficient in front of Cl2 .

There is 2 Cl on left side and 4 Cl on right side. By writing 2 in front of Cl2 , we can make Cl equal on both the side.


SO2 + 2 Cl2 –> SOCl2 + Cl2O

Reactant / Product


S 1 / 1


O 2 / 1 + 1 = 2


Cl 4 / 2 + 2 = 4



Coefficients are the numbers in front of each formula. If no number is shown, a 1 is understood.


1 SO2 + 2 Cl2 –> 1 SOCl2 + 1 Cl2O


So after balancing the reaction, the coefficient of SOCl2 is 1.




Have a look at the picture attached.

Thank you. :)